how can i write six times the difference of a number and 9 is 54 as an equation?.
Answer by Leaf W.(135)

x = "a number"
x - 9 = "the difference of a number and nine"
6(x - 9) = "six times the difference of a number and 9"
==> Therefore, the equation is 6(x - 9) = 54